# Concomitant Entanglement and Control Criticality Driven by Collective Measurements

### v1

@@ III.2 Quantum Control Transition
Figure 6: Order parameter (quantum model). Shown at time $t=L$ as a function of control rate $p$ for the quantum model. Inset: Finite-size scaling collapse indicates $p_{c}=0.46(3)$, $\beta=0.28(4)$, $\nu=2.0(2)$. $A\approx 21.7$ and $\alpha\approx 1.46$ provide the leading corrections to finite-size scaling.
@@
In Fig. 6, we plot the order parameter $\overline{\braket{H_{ZZ}(t=L)}}$ averaged over $10^{4}$ circuit realizations for system sizes $L=16,\dots,32$ and $p=0.1,\dots,0.7$. The results are qualitatively similar to the small-system results for the classical model shown in Fig. 4: the order parameter grows with $L$ at small $p$ and shrinks with $L$ at large $p$, with a finite-size crossing near $p=0.5$. However, our analysis of the small-system classical data demonstrated that this crossing was a red herring, with the large-system data indicating a transition well below the small-system crossing point with the order parameter approaching zero rather than remaining finite. With this in mind, we perform a finite-size scaling analysis of the quantum data assuming a critical scaling form given in Eq. (12) for the order parameter data, which can take into account the expected substantial finite-size effects. The data collapse obtained from a 5-parameter $\chi^{2}$ analysis is shown in the inset of Fig. 6. The analysis yields an estimate of the transition location $p_{c}=0.46(3)$, order-parameter exponent $\beta=0.28(4)$, and correlation-length exponent $\nu=2.0(2)$. Finite-size effects are accounted for by $A\approx 21.7$ and $\alpha\approx 1.46$, leading to finite-size corrections ranging from $AL^{-\alpha}\approx 0.38$ ($L=16$) to $0.14$ ($L=32$), which is on the same order as those obtained for the small-system classical data. The transition location $p_{c}$ and correlation-length exponent $\nu$ agree within error bars with those obtained for the classical transition, while the order parameter exponent $\beta$ is substantially larger (by about a factor of 2) for the quantum transition.
@@
To extract the dynamical critical exponent $z$, in Fig. 7 we show the dynamics of the order parameter for various system sizes at $p=0.46$, near the control transition. Owing to the substantial finite size effects observed in the late-time value of the order parameter considered above, we allow for a system-size dependent but time independent shift of the order parameter dynamics such that $L^{\beta/\nu}\overline{\braket{H_{ZZ}(t)}}+BL^{-\gamma}=f(t/L^{z})$ is our scaling function [which can be obtained by Taylor expanding Eq. (12)]. Using the value $\beta/\nu=0.14$ obtained from Fig. 6, we perform a 3-parameter $\chi^{2}$ analysis that yields $z=0.84(4)$ (see inset). Finite-size effects are accounted for by $B\approx 18.0$ and $\gamma\approx 2.11$, corresponding to finite size corrections ranging from $BL^{-\gamma}\approx 0.03$ ($L=20$) to $0.01$ ($L=32$). While these corrections are about an order of magnitude smaller than those obtained for the late-time order parameter data, they should be compared to the scale of the vertical axis in Fig. 7, which is also an order of magnitude smaller than that of the horizontal axes in Figs. 6 and 4. Thus the finite size corrections to our scaling analyses are broadly consistent. Moreover, the value $z=0.84(4)$ matches the value $z=0.84(9)$ obtained for the classical transition in Fig. 5. This further indicates that the “long-range” nature of the classical transition persists in the quantum limit.
@@
Figure 7: Dynamics of the order parameter (quantum model). Shown at $p=0.46$, near the control transition. The inset shows a scaling collapse using the values of $\beta$ and $\nu$ extracted from Fig. 6. The collapse yields an estimated dynamical exponent $z=0.84(4)$. $B\approx 18.0$ and $\gamma\approx 2.11$ provide the leading corrections to finite-size scaling.
@@ IV Entanglement Transition
where $\Gamma(p,L)$ is a decay rate and the scaling of $S_{\infty}(p,L)$ encodes the properties of the entanglement phase. Further details on this decay, and a discussion of how to remove it entirely, are discussed in Appendix B. We expect that these signatures are present in previously studied absorbing-state transitions [15, 16, 17], although the effects may be small if large systems are being studied with, e.g., Clifford or free-fermion methods. It is also possible that the probability of disentangling the system by becoming accidentally trapped on the absorbing state is larger in the presence of global measurements like those employed in our control protocol. Quantifying these differences could help elucidate controllability within the chaotic (uncontrolled) regime; we leave this for future work.
An important observation made in Appendix B is that the decay rate $\Gamma(p,L)$ is approximately $L$-independent close to the transition and deep in the area-law phase. Thus, evaluating the entanglement entropy at a time $t$ of order $L$ allows us to probe the criticality via the entanglement scaling encoded in $S_{\infty}(p,L)$ in Eq. (13), although some residual noise in the data due to these systematics is still to be expected.
@@ IV.1 Entanglement Entropy Scaling
Fig. 9 shows the TMI at time $t=L$ as a function of $p$ for $L=16,\dots,28$. It displays the expected behavior at large and small $p$ and appears to manifest several finite-size crossings between $p\approx 0.45$ and $0.55$, where the largest and smallest pair of consecutive system sizes, respectively, exhibit crossings. The lack of a single clear crossing is likely due to the finite-size effects discussed in Sec. III and Appendix B. To mitigate these finite-size effects, we perform a scaling collapse using the same ansatz adopted in Fig. 6, with the result shown in the inset. This collapse yields an estimated critical point $p_{c}=0.44(2)$ and correlation-length exponent $\nu=1.6(2)$, consistent within error bars with the corresponding quantities extracted for the control transition in Fig. 4. The best collapse is achieved by applying the finite size scaling ansatz
$\overline{\mathcal{I}_{3}(t/L^{z}\to\infty)}_{p=p_{c}}\sim g\left[L^{1/\nu}(p-p_{c})+A_{Q}L^{-\alpha}\right]$ (15)
@@ IV.2 Purification Transition
To further pinpoint the purification transition and estimate the correlation-length critical exponent $\nu$, we use the system-size scaling of $\overline{S_{\rm anc}}$ evaluated at a time of order $L^{z}$. Measuring this quantity as a function of $p$ for various $L$, we expect to see curves for different $L$ cross at $p_{c}$, with $\overline{S_{\rm anc}}$ increasing (decreasing) as a function of $L$ below (above) this value. This is shown in Fig. 11, which plots $\overline{S_{\rm anc}}$ at time $t=0.7L^{0.88}$ (we confirmed that the results described below do not depend strongly on the choice of the prefactor $0.7$—any prefactor $\gtrsim 0.6$ works well as long as the value of $\overline{S_{\rm anc}}$ is not too small). Note that, since this protocol requires obtaining $\overline{S_{\rm anc}}$ at noninteger times, we used a linear interpolation to estimate its value and determine the associated error bars from those of the neighboring data points. A scaling collapse assuming a scaling function $f[L^{1/\nu}(p-p_{c})]$ is then performed (see inset), yielding the estimates $p_{c}=0.41(2)$ and $\nu=1.7(2)$. These values serve as independent estimates of the location and critical properties of the entanglement transition, and are consistent with those obtained from the TMI data in Fig. 9. Notably, however, the finite size effects are much less pronounced for the ancilla entropy, making it perhaps the most reliable witness of the entanglement transition. In fact, carrying out the collapse with a scaling ansatz of the form (15) yields finite size corrections that are zero to machine precision, giving the same $p_{c}$, $\nu$, and error bars.
V Discussion and Outlook
@@
We have studied the impact of collective measurements on quantum control and entanglement dynamics in a model inspired by Rydberg atom quantum simulators and by the classical notion of probabilistic control of chaos. The critical points and critical exponents extracted from the various metrics considered in our study are collected for reference in Table 1. With the aid of a numerically tractable classical limit, we identified a control transition driven by collective measurements and local adaptive correction that features a dynamical critical exponent $z<1$, indicating “superballistic” spreading of correlations at criticality. The correlation-length and dynamical exponents of the transition obtained in the classical and quantum limits of the model are in agreement. In the quantum limit of the model, we also establish the existence of an entanglement transition witnessed by two independent measures, $\overline{\mathcal{I}_{3}}$ and $\overline{S_{\rm anc}}$. The estimates of the entanglement critical point extracted from both measures are indistinguishable from the estimated location of the control transition to within the resolution of our finite-size numerics. Furthermore, both entanglement metrics estimate a correlation length critical exponent $\nu$ between $1.4$ and $1.9$, which is consistent with the value $2.0(2)$ estimated from the control order parameter. Finally, the dynamical critical exponent $z<1$ estimated from the dynamics of $S_{\rm anc}$ is consistent with the value estimated from the dynamics of the control order parameter $H_{ZZ}$ in both the quantum and classical limits.
@@
$\overline{\braket{H_{ZZ}}}$ (classical) 0.49(1) 2.1(2) 0.16(2) 0.84(9)
$\overline{\braket{H_{ZZ}}}$ (quantum) 0.46(3) 2.0(2) 0.28(4) 0.84(4)
@@
Table 1: Transition points and critical exponents. Extracted from the various metrics considered in this paper.
@@
It is interesting to consider the implications of our results for the topology of the control-entanglement phase diagram—in particular, do the control and entanglement transitions coincide, such that the control criticality servers as a reliable witness of the entanglement transition? Two scenarios present themselves. First, the control and entanglement transitions indeed coincide, and their critical properties match. Second, the control and entanglement transitions occur very close to one another but do not coincide. The first scenario is consistent with the observation that control and entanglement transitions can coincide in the presence of a nonlocal control protocol [7, 11, 10]. Furthermore, in cases where the two transitions occur nearby but do not coincide, typically it is still possible to deduce that the transitions are different by observing a difference in critical properties. For example, in Ref. [16] the entanglement and absorbing-state transitions can be distinguished by their value of $z$ even when the transitions are brought close together. Typically, in this scenario, the entanglement transition belongs to the universality class of the MIPT [56] with $z=1$ and $\nu\approx 1.3$, while universality of the control transition is often governed by classical physics, e.g. a random walk in Ref. [7] and the directed-percolation or parity-conserving universality classes in Refs. [16, 17], all of which feature $z>1$. In the present case, the entanglement and control transitions have consistent dynamical critical exponents, and the correlation length critical exponents $\nu$ for the two transitions are also not distinguishable within the resolution of our numerics. This further supports the hypothesis that the transitions coincide.
@@
If the transitions do not coincide, then the entanglement transition manifests critical exponents associated with a nonstandard MIPT. Given the dynamical exponent $z<1$, it is natural to look to the MIPT with long-range scrambling gates studied in Ref. [34]. There, two-qubit Clifford gates are applied between qubits a distance $r$ apart with probability $P(r)\sim 1/r^{\alpha}$, with $\alpha=0$ corresponding to all-to-all interactions. Ref. [34] found an MIPT with exponents that vary continuously as a function of $\alpha$, with standard short-range MIPT criticality emerging for $\alpha\gtrsim 3.0$. Intriguingly, an MIPT with $z\approx 0.88$ (matching the dynamical exponent of the entanglement transition studied here) occurs when $\alpha\approx 2.9$. The corresponding correlation length critical exponent $\nu\approx 1.4$, which is within error bars of the correlation length exponent $\nu=1.6(2)$ determined from the TMI, but just outside error bars of the value $\nu=1.7(2)$ determined from the ancilla entropy. However, this comparison should be made with caution as MIPTs in Clifford circuits are known to exhibit distinct universality relative to, e.g., Haar random circuits [56]. Moreover, even if the transitions ultimately do not coincide, the fact that their dynamical exponents are so similar seems to suggest that they are intertwined in some way.
@@
This points to one worthwhile direction for future work, which is to study how the standard MIPT is affected by the inclusion of magnetization measurements. This would be useful in establishing a baseline to which the critical properties of the entanglement transition studied here can be compared. Another particularly interesting question is how to understand the emergence of a dynamical exponent $z<1$ in the entanglement transition. For example, is it possible to connect the critical dynamics at the MIPT to a model with explicit long-range interactions? Further, the question about the nature of the volume-law phase in the presence of magnetization measurements needs further exploration. For example, in monitored random circuits the entanglement in the volume-law phase acquires a subleading system-size dependent correction related to the roughness exponent of a directed polymer in a random environment [65]. Does a new exponent emerge in the presence of long-range measurements? Finally, it is worthwhile to consider the utility of collective measurements for quantum-technology applications including state preparation, sensing, and quantum error correction.

### v2

@@ III.2 Quantum Control Transition
Figure 6: Order parameter (quantum model). Shown at time $t=L^{0.85}$ as a function of control rate $p$ for the quantum model. Inset: Finite-size scaling collapse indicates $p_{c}=0.46(5)$, $\beta=0.32(4)$, $\nu=2.05(25)$. $A\approx 28.2$ and $\alpha\approx 1.54$ provide the leading corrections to finite-size scaling.
@@
Our analysis of the quantum control transition follows that of the classical transition presented in the previous subsection. In Fig. 6, we plot the order parameter $\overline{\braket{H_{ZZ}(t=L^{0.85})}}$ averaged over $10^{4}$ circuit realizations for system sizes $L=16,\dots,32$ and $p=0.2,\dots,0.7$. The exponent $0.85$ is chosen to achieve the best collapse, and a direct analysis of the dynamics near the critical point (Fig. 7) yields an estimated dynamical exponent $z=0.85(5)$ (see discussion below). The results are qualitatively similar to the small-system results for the classical model shown in Fig. 4: the order parameter grows with $L$ at small $p$ and shrinks with $L$ at large $p$, with a finite-size crossing near $p=0.5$. However, our analysis of the small-system classical data demonstrated that this crossing was a red herring, with the large-system data indicating a transition well below the small-system crossing point with the order parameter approaching zero rather than remaining finite. With this in mind, we perform a finite-size scaling analysis of the quantum data assuming a critical scaling form given in Eq. (12) for the order parameter data, which can take into account the expected substantial finite-size effects. The data collapse obtained from a 5-parameter $\chi^{2}$ analysis is shown in the inset of Fig. 6. The analysis yields an estimate of the transition location $p_{c}=0.46(5)$, order-parameter exponent $\beta=0.32(4)$, and correlation-length exponent $\nu=2.05(25)$. Finite-size effects are accounted for by $A\approx 28.2$ and $\alpha\approx 1.54$, leading to finite-size corrections ranging from $AL^{-\alpha}\approx 0.39$ ($L=16$) to $0.14$ ($L=32$), which is on the same order as those obtained for the small-system classical data. The transition location $p_{c}$ and correlation-length exponent $\nu$ agree within error bars with those obtained for the classical transition, while the order parameter exponent $\beta$ is substantially larger (by about a factor of 2) for the quantum transition.
@@
To extract the dynamical critical exponent $z$, in Fig. 7 we show the dynamics of the order parameter for various system sizes at $p=0.46$, near the control transition. Owing to the substantial finite size effects observed in the late-time value of the order parameter considered above, we allow for a system-size dependent but time independent shift of the order parameter dynamics such that $L^{\beta/\nu}\overline{\braket{H_{ZZ}(t)}}+BL^{-\gamma}=f(t/L^{z})$ is our scaling function [which can be obtained by Taylor expanding Eq. (12)]. Using the value $\beta/\nu=0.157$ obtained from Fig. 6, we perform a 3-parameter $\chi^{2}$ analysis that yields $z=0.85(5)$ (see inset). Finite-size effects are accounted for by $B\approx 8.06$ and $\gamma\approx 1.76$, corresponding to finite size corrections ranging from $BL^{-\gamma}\approx 0.04$ ($L=20$) to $0.02$ ($L=32$). While these corrections are about an order of magnitude smaller than those obtained for the late-time order parameter data, they should be compared to the scale of the vertical axis in Fig. 7, which is also an order of magnitude smaller than that of the horizontal axes in Figs. 6 and 4. Thus the finite size corrections to our scaling analyses are broadly consistent. Moreover, the value $z=0.85(5)$ matches the value $z=0.8(1)$ obtained for the classical transition in Fig. 5. This further indicates that the “long-range” nature of the classical transition persists in the quantum limit.
@@
Figure 7: Dynamics of the order parameter (quantum model). Shown at $p=0.46$, near the control transition. The inset shows a scaling collapse using the values of $\beta$ and $\nu$ extracted from Fig. 6. The collapse yields an estimated dynamical exponent $z=0.85(5)$. $B\approx 8.06$ and $\gamma\approx 1.76$ provide the leading corrections to finite-size scaling.
@@ IV Entanglement Transition
where $\Gamma(p,L)$ is a decay rate and the scaling of $S_{\infty}(p,L)$ encodes the properties of the entanglement phase. Further details on this decay, and a discussion of how to remove it entirely, are discussed in Appendix C. We expect that these signatures are present in previously studied absorbing-state transitions [15, 16, 17], although the effects may be small if large systems are being studied with, e.g., Clifford or free-fermion methods. It is also possible that the probability of disentangling the system by becoming accidentally trapped on the absorbing state is larger in the presence of global measurements like those employed in our control protocol. Quantifying these differences could help elucidate controllability within the chaotic (uncontrolled) regime; we leave this for future work.
An important observation made in Appendix C is that the decay rate $\Gamma(p,L)$ is approximately $L$-independent close to the transition and deep in the area-law phase. Thus, evaluating the entanglement entropy at a time $t$ of order $L$ allows us to probe the criticality via the entanglement scaling encoded in $S_{\infty}(p,L)$ in Eq. (13), although some residual noise in the data due to these systematics is still to be expected.
@@ IV.1 Entanglement Entropy Scaling
Fig. 9 shows the TMI at time $t=L$ as a function of $p$ for $L=16,\dots,28$. It displays the expected behavior at large and small $p$ and appears to manifest several finite-size crossings between $p\approx 0.45$ and $0.55$, where the largest and smallest pair of consecutive system sizes, respectively, exhibit crossings. The lack of a single clear crossing is likely due to the finite-size effects discussed in Sec. III and Appendix C. To mitigate these finite-size effects, we perform a scaling collapse using the same ansatz adopted in Fig. 6, with the result shown in the inset. This collapse yields an estimated critical point $p_{c}=0.44(2)$ and correlation-length exponent $\nu=1.6(2)$, consistent within error bars with the corresponding quantities extracted for the control transition in Fig. 4. The best collapse is achieved by applying the finite size scaling ansatz
$\overline{\mathcal{I}_{3}(t/L^{z}\to\infty)}_{p>p_{c}}\sim g\left[L^{1/\nu}(p-p_{c})+A_{Q}L^{-\alpha}\right]$ (15)
@@ IV.2 Purification Transition
To further pinpoint the purification transition and estimate the correlation-length critical exponent $\nu$, we use the system-size scaling of $\overline{S_{\rm anc}}$ evaluated at a time of order $L^{z}$. Measuring this quantity as a function of $p$ for various $L$, we expect to see curves for different $L$ cross at $p_{c}$, with $\overline{S_{\rm anc}}$ increasing (decreasing) as a function of $L$ below (above) this value. This is shown in Fig. 11, which plots $\overline{S_{\rm anc}}$ at time $t=0.7L^{0.88}$ (we confirmed that the results described below do not depend strongly on the choice of the prefactor $0.7$—any prefactor $\gtrsim 0.6$ works well as long as the value of $\overline{S_{\rm anc}}$ is not too small). A scaling collapse assuming a scaling function $f[L^{1/\nu}(p-p_{c})]$ is then performed (see inset), yielding the estimates $p_{c}=0.41(2)$ and $\nu=1.7(2)$. These values serve as independent estimates of the location and critical properties of the entanglement transition, and are consistent with those obtained from the TMI data in Fig. 9. Notably, however, the finite size effects are much less pronounced for the ancilla entropy, making it perhaps the most reliable witness of the entanglement transition. In fact, carrying out the collapse with a scaling ansatz of the form (15) yields finite size corrections that are zero to machine precision, giving the same $p_{c}$, $\nu$, and error bars.
V Exploring the Phase Diagram
So far, we have focused on the control and entanglement transitions as a function of the control rate $p$. However, the properties and locations of the two transitions also depend on the parameter $q$ (fixed at $0.2$ in Secs. III and IV), which sets the fraction of sites on which local domain wall corrections are performed. Intuitively, we expect $p_{c}$ to drift upwards with decreasing $q$. At $q=1$, when all domain walls are measured and corrected, the system immediately disentangles and becomes controlled in an $O(1)$ time, so $p_{c}=0$. At $q=0$, when no corrections are performed, the system never becomes controlled and remains volume-law entangled for any $p$, so $p_{c}\to\infty$. As $q$ is increased, we expect the dynamical exponent $z$ to flow downwards and the correlation length exponent $\nu$ to flow upwards. As more domain-wall measurements and corrections are performed, we expect control to occur more rapidly in time, and this should be reflected in the timescale for control at criticality via a decrease in $z$. Moreover, since the domain-wall measurements and corrections are conditioned on the same collective measurement outcomes, we expect them to generate correlations in space that should be reflected in the divergence of the correlation length by an increase in $\nu$.
Figure 12: Phase diagram (quantum model). The locations of the entanglement ($\overline{S_{\rm anc}}$) and control ($\overline{\braket{H_{ZZ}}}$) transitions, along with their error bars from $\chi^{2}$ analysis, are shown in red and purple, respectively. Here, $p$ is the probability that the control circuit is applied at each time step, and $q$ is the fraction of sites undergoing a local domain wall correction when the control is applied. The shaded region represents the volume-law uncontrolled phase. The dashed line at $q=0.2$ represents the line cut considered in previous sections.
Figure 13: Evolution of critical exponents (quantum model). Correlation length exponents $\nu$ (red) and dynamical exponents $z$ (blue) for the entanglement (darker circles) and control transitions (lighter diamonds) are plotted along with their error bars as a function of the local correction density $q$. The dashed line at $q=0.2$ represents the transition studied in previous sections.
Figs. 12 and 13 largely support this intuition. In Fig. 12 we plot a portion of the phase diagram as a function of $p$ and $q$, focusing on $\overline{S_{\rm anc}}$ as a reliable probe of the entanglement transition. As expected, we see that $p_{c}$ (red points) drifts to larger values as $q$ decreases, implying that the control circuit must be applied more frequently in order to drive the entanglement transition. To test whether the locations of the entanglement and control transitions continue to coincide as $q$ is varied, we also plot the values of $p_{c}$ extracted from $\overline{\braket{H_{ZZ}}}$ (purple points) for $q=0.3$ and $0.08$, in addition to the value obtained for $q=0.2$ (dashed line) in the previous section. We find that the transitions coincide up to the resolution of our small-size numerics for $q=0.3$ and $0.2$, but that the transitions appear to separate for $q=0.08$.
In Fig. 13 we plot the critical exponents $\nu$ (red) and $z$ (blue) along the phase boundary as $q$ is varied for both the entanglement (darker circles) and control transitions (lighter diamonds). For both transitions, we see the expected trend of increasing $\nu$ and decreasing $z$ as $q$ increases. We also see that the dynamical exponents of the two transitions remain very close to one another, and well within error bars, as $q$ varies. Notably, for small $q\lesssim 0.16$, the dynamical exponent $z$ becomes greater than $1$, signaling that the transitions are recovering a “short-range” nature with critical correlations spreading ballistically or slower. The values of $\nu$ for the entanglement and control transitions coincide up to the resolution of our numerics for $q=0.2$ and $0.3$ where the transitions are “long-range” with $z<1$, while they differ by several error bars at $q=0.08$ where the transitions have $z\approx 1.1$. Making a more conclusive statement on the respective values of $p_{c}$, $z$, and $\nu$ for the two transitions requires system sizes beyond those accessible to our exact numerics. This may be achievable by applying matrix product state techniques from the controlled side of the transition, which is a worthwhile subject for future work.
VI Discussion and Outlook
@@
We have studied the impact of collective measurements on quantum control and entanglement dynamics in a model inspired by Rydberg atom quantum simulators and by the classical notion of probabilistic control of chaos. The critical points and critical exponents extracted from the various metrics considered in our study at $q=0.2$ are collected for reference in Table 1. With the aid of a numerically tractable classical limit, we identified a control transition driven by collective measurements and local adaptive correction that features a dynamical critical exponent $z<1$, indicating “superballistic” spreading of correlations at criticality. The correlation-length and dynamical exponents of the transition obtained in the classical and quantum limits of the model are in agreement. In the quantum limit of the model, we also establish the existence of an entanglement transition witnessed by two independent measures, $\overline{\mathcal{I}_{3}}$ and $\overline{S_{\rm anc}}$. The estimates of the entanglement critical point extracted from both measures are indistinguishable from the estimated location of the control transition to within the resolution of our finite-size numerics. Furthermore, both entanglement metrics estimate a correlation length critical exponent $\nu$ between $1.4$ and $1.9$, which is consistent with the value $2.05(25)$ estimated from the control order parameter. Finally, the dynamical critical exponent $z<1$ estimated from the dynamics of $S_{\rm anc}$ is consistent with the value estimated from the dynamics of the control order parameter $H_{ZZ}$ in both the quantum and classical limits.
Upon varying $q$, we find that the entanglement and control transitions retain a “long-range” nature with $z<1$ down to $q\sim 0.16$, below which they revert to a “short-range” nature with $z\gtrsim 1$. In the long-range regime with $z<1$, the locations and critical exponents of the control and entanglement transitions coincide up to the resolution of our numerics. In the short-range regime, we find evidence that the transitions begin to take on distinct locations and critical properties, including different values of the correlation length exponent $\nu$. We conjecture that the long-range character of the transition for $q\gtrsim 0.16$ is essential to the apparent coincidence of the transitions. Our model allows access to kinetically constrained MIPTs as $q\to 0$ and infinitely fast transitions with $z\to 0$ as $q\to 1$. The ability to explore these regimes as a function of $p$ and $q$ is an exciting outcome of our study.
@@
$\overline{\braket{H_{ZZ}}}$ (classical) 0.49(1) 2.3(2) 0.17(2) 0.8(1)
$\overline{\braket{H_{ZZ}}}$ (quantum) 0.46(5) 2.05(25) 0.32(4) 0.85(5)
@@
Table 1: Transition points and critical exponents at $\bm{q=0.2}$. Extracted from the various metrics considered in this paper.
@@
It is interesting to consider the implications of our results for the topology of the control-entanglement phase diagram—in particular, do the control and entanglement transitions coincide, such that the control criticality serves as a reliable witness of the entanglement transition? Two scenarios present themselves. First, the control and entanglement transitions indeed coincide, and their critical properties match. Second, the control and entanglement transitions occur very close to one another but do not coincide. The first scenario is consistent with the observation that control and entanglement transitions can coincide in the presence of a nonlocal control protocol [7, 11, 10]. Furthermore, in cases where the two transitions occur nearby but do not coincide, typically it is still possible to deduce that the transitions are different by observing a difference in critical properties. For example, in Ref. [16] the entanglement and absorbing-state transitions can be distinguished by their value of $z$ even when the transitions are brought close together. Typically, in this scenario, the entanglement transition belongs to the universality class of the MIPT [62] with $z=1$ and $\nu\approx 1.3$, while universality of the control transition is often governed by classical physics, e.g. a random walk in Ref. [7] and the directed-percolation or parity-conserving universality classes in Refs. [16, 17], all of which feature $z>1$. In the present case, the entanglement and control transitions have consistent dynamical critical exponents. Moreover, the correlation length critical exponents $\nu$ for the two transitions are also not distinguishable within the resolution of our numerics, except at small local correction density $q$. This further supports the hypothesis that the transitions coincide for sufficiently large $q\gtrsim 0.16$.
@@
If the transitions do not coincide, then the entanglement transition manifests critical exponents associated with a nonstandard MIPT. Given the dynamical exponent $z<1$, it is natural to look to the MIPT with long-range scrambling gates studied in Ref. [34]. There, two-qubit Clifford gates are applied between qubits a distance $r$ apart with probability $P(r)\sim 1/r^{\alpha}$, with $\alpha=0$ corresponding to all-to-all interactions. Ref. [34] found an MIPT with exponents that vary continuously as a function of $\alpha$, with standard short-range MIPT criticality emerging for $\alpha\gtrsim 3.0$. Intriguingly, an MIPT with $z\approx 0.88$ (matching the dynamical exponent of the entanglement transition studied here at $q=0.2$) occurs when $\alpha\approx 2.9$. The corresponding correlation length critical exponent $\nu\approx 1.4$, which is within error bars of the correlation length exponent $\nu=1.6(2)$ determined from the TMI, but just outside error bars of the value $\nu=1.7(2)$ determined from the ancilla entropy. However, this comparison should be made with caution as MIPTs in Clifford circuits are known to exhibit distinct universality relative to, e.g., Haar random circuits [62]. Moreover, even if the transitions ultimately do not coincide, the fact that their dynamical exponents are so similar seems to suggest that they are intertwined in some way.
@@
This points to one worthwhile direction for future work, which is to study how the standard MIPT is affected by the inclusion of magnetization measurements. This would be useful in establishing a baseline to which the critical properties of the entanglement transition studied here can be compared. Another particularly interesting question is how to understand the emergence of a dynamical exponent $z<1$ in the entanglement transition. This dynamical exponent indicates that collective measurements are enhancing teleportation events similar to how rare regions can within a local projective measurement context [72]. For example, is it possible to connect the critical dynamics at the MIPT to a model with explicit long-range interactions? Further, the question about the nature of the volume-law phase in the presence of magnetization measurements needs further exploration. For example, in monitored random circuits the entanglement in the volume-law phase acquires a subleading system-size dependent correction related to the roughness exponent of a directed polymer in a random environment [73]. Does a new exponent emerge in the presence of long-range measurements?
Our results demonstrate how collective measurements in adaptive quantum dynamics can be harnessed to robustly control onto dynamical states, which is relevant for the development of efficient quantum algorithms and error correction protocols in NISQ devices.